# MMVP: A Multimodal MoCap Dataset with Vision and Pressure Sensors

### v1

@@ 4.2 Pose and Translation Optimization
Since CLIFF [31] has a good perception of global rotation, we utilize CLIFF to obtain initial pose $\tilde{\boldsymbol{\theta}}$, shape $\tilde{\boldsymbol{\beta}}$, and rotation $\tilde{\boldsymbol{R}}$. We do not optimize global rotation and instead adopt the results from CLIFF. It is important to note that a single set of image sequences corresponds to a specific individual. Therefore, we fix an average of $\tilde{\boldsymbol{\beta}}$ for the same person across the entire image sequence. To simplify the notation, we use $\boldsymbol{J}(\boldsymbol{\theta},\boldsymbol{T})$ to represent $\boldsymbol{J}(\boldsymbol{\theta},\tilde{\boldsymbol{\beta}},\tilde{\boldsymbol{R}},\boldsymbol{T})$.
@@
Contact joint loss. In order to recover the positional relationship between the human and the ground, we utilize the off-the-shelf depth estimation model ZoeDepth [2] to obtain depth map for each RGB image, and use the corresponding point cloud $\boldsymbol{P}$ to construct the 3D contact trace on the ground. Therefore, the spatial translation constraint term can be defined by penalizing the 3D distance between the SMPL foot joint $\boldsymbol{J}(\boldsymbol{\theta},\bar{\boldsymbol{\beta}})_{foot}$ and the contact point cloud $\boldsymbol{P}(\mathcal{J}^{2d}_{foot})$:
@@
Foot consistency loss. Meanwhile, we assume that the foot joints position of two consecutive frames are consistently close during contact. At the same time, the jitter of 2D keypoints $\mathcal{J}^{2d}_{foot}$ will cause the corresponding contact joint on the ground to shift significantly. In order to solve the above problems, we define a temporal constraint term as,
@@ 5.1 Comparison of Ground Truth Registration
We compare our RGBD-P fitting method with two existing RGBD methods: PROX [14] and LEMO [61]. Both of these methods take human-scene contact into consideration in pose fitting. We employ the 3D depth error $E_{3d}$ to quantify the alignment between human models and depth observations. To evaluate the interaction between the model and the scene, we calculate the model contact by setting a unified threshold and compare the mean per foot vertex error (MPFVE), F1 score, and IOU with the ground truth contact. Tab. 2 demonstrates the effectiveness of proposed method.
@@
Table 2: Evaluation of pose and shape registration on MMVP. M.: MPFVE.
@@ 5.2 Comparison of Foot Contact Estimation
We apply precision, recall, F1 score, and IOU as evaluation metrics and conduct a quantitative comparison with BSTRO [18], BSTRO[FT], and DECO [49] on MMVP-test (Tab. 3). BSTRO[FT] represents BSTRO after fine-tuning on the MMVP training dataset. From Tab. 3, it is evident that our method achieves the best performance. It is worth mentioning that BSTRO and DECO, due to contact annotations in their training dataset (illustrated in Fig. 2), have a tendency to predict feet as contact in the majority of cases. As a result, their recall values are very high, while other metrics are undesirable. It is challenging to perceive the foot contact solely based on a single-frame image. However, by incorporating temporal motion information, we believe this is the reason why proposed method outperforms the fine-tuned BSTRO.
To validate the generalization, we also compare our method with BSTRO[FT] on public datasets. Due to the inaccurate contact labels in other datasets, we only conduct qualitative comparisons (Fig. 5). It can be observed that proposed method not only achieves more accurate contact predictions but also estimates reasonable foot pressure distributions. For more details, please refer to the supplementary material.
Figure 6: Qualitative comparison of 3D translation estimation in Side-Slip (Left), Standing-Long-Jump (Middle) and Throwing-Medicine-Ball (Right) case. We estimate 3D translation stability and compare our results (Pink) with CLIFF [31] (Blue) and TRACE [48] (Green). The lowest row shows the changes of acceleration over time for each corresponding case.
@@ 5.3 Pose and Translation Optimization Results
To assess the accuracy of our 3D pose estimation, we employ various metrics: MPJPE (Mean Per Joint Position Error), PMPJPE (Procrustes-aligned Mean Per Joint Position Error), PVE (Per Vertex Error) and PVE (Feet) which specifically focuses on the error of SMPL mesh vertices corresponding to the soles of the feet. For evaluating translation estimation, we utilize the offset distance of the pelvis point relative to the starting position, denoted as Traj.
Quantitative results. We conduct a comparative analysis of our method against CLIFF [31], TRACE [48], and SMPLer-X [5] within our dataset. As depicted in Tab. 5, our method exhibits superior performance across all metrics on MMVP. Notably, our method excels in Traj, indicating a significant improvement in acquiring more accurate and robust translation by leveraging foot contact for SMPL fitting.
Qualitative results. We compare 3D translation between our method, CLIFF [31] and TRACE [48] on our dataset (Fig. 6). In the Side-Slip case, both CLIFF and TRACE exhibit drift in the depth direction in their respective 3D translation estimation, whereas our method demonstrates a consistent and stable translation in the depth direction. Similarly, in the Standing-Long-Jump and Throwing-Medicine-Ball case, the 3D translation estimated by CLIFF and TRACE manifests drift in the depth direction during the preparation stage. In contrast, our method avoids any depth-direction translation. At the bottom, we show the averaged absolute acceleration of feet joints. Our method achieves the smallest value, indicating superior foot stability.
5.4 Ablation
To investigate the impacts of $L_{3d}$ and $L_{t}$, we conduct an ablation experiment on our dataset. As shown in Tab. 5, the omission of $L_{t}$ results in a marginal increase in all error metrics. This observation indicates that the temporal foot contact signals serve not only to improve the accuracy of translation but also improve the accuracy of pose. Without $L_{3d}$, MPJPE, PA-MPJPE and PVE have dropped slightly, but PVE (Feet) and Traj has increased a lot. This is because when the 3D constraints of the feet and the ground are introduced, it will affect the 2D constraints to a certain extent and cause the accuracy of pose to drop a little, but in exchange for the accuracy of the soles of the feet vertices and global translation. In order to attain enhanced translation, it is necessary to make a trade-off by sacrificing a certain degree of pose accuracy. Additionally, getting a better translation requires finding a balance between 2D and 3D constraints.
@@ 6 Conclusion
Limitations and future work: Our method is mainly designed for videos from static viewpoints, so it remains challenging to handle moving camera configurations. The existing MMVP have relatively homogeneous scenes, making it difficult for current visual networks to achieve scene generalization on this dataset. In the future, we plan to collect data from various outdoor scenes to expand the dataset. Additionally, it would be an intriguing research endeavor to learn dynamic priors solely from pressure sequences.
Acknowledgements: This work was supported by Xtep Science Lab.
@@ 7 Dataset Details
Statistics:The dataset consists of a total of 16 subjects, with 11 males and 5 females. Key statistics (gender, age, height and weight) of the subjects are shown in Fig. 7. Synchronized RGBD frames, pressure data, and registrations are more than 44k frames. The dataset consists of high-speed and large-range movements, including skipping rope, long jump, ball throwing, side stepping, running, and dancing.
@@ 8.1 Foot Contact Temporal Loss
We design four foot planes derived from SMPL template mesh, as shown in Fig. 8. By projecting the foot vertices of template SMPL model onto the ground, we acquire the anterior and posterior foot planes. The points on foot planes, which are associated with SMPL foot vertices, also can be controlled by SMPL parameters.
@@
8.2 RGBD-P Fitting Details
@@ 8.3 RGBD-P fitting Comparison Details
We compare our results with PROX  [^3] https://github.com/mohamedhassanmus/prox and LEMO  [^4] https://github.com/sanweiliti/LEMO on MMVP Dataset. All configs are set as their default values. Compared to PROX, our method takes account body shape difference between adults and children, ensuring shape parameters consistence in one motion sequence. PROX sets a distance threshold between foot and ground, which may result in sudden motion. To avoid this, our method leverages dense foot-ground distance as loss function in optimization. Based on the fitting results of PROX, LEMO considers temporal motion changing and contact consistence, however, it neglects the alignment with depth data, which can lead to difficulties in accurately estimating global position. For more comparison results, please refer to our supplemental video.
@@ 9.1 FPP-Net
Network details. We follow PhysCap  [^5] https://github.com/soshishimada/PhysCap_demo_release and Jesse et al. [44] to construct our network. The details are shown in Fig. 9. The numbers next to ”Conv.1d” represent the input dimension, output dimension, and kernel size. The numbers next to ”Linear” represent the input dimension and output dimension. The outputs of FPP-Net are marked in purple. We train our network for 20 hours (900 epochs). The initial learning rate is $10^{-4}$. The optimizer is Adam.
Dataset split. We split 16 subjects in MMVP into 13, 3 for training and testing, respectively.
@@ 9.4 More results on in-the-wild datasets
Table 5: Evaluation on 3DPW and EMDB.
We compared the results on in-the-wild datasets, 3DPW [53] and EMDB [24]. Our method is mainly designed for videos from static viewpoints, so it remains challenging to handle moving camera configurations. Therefore, we selected sequences with relatively static camera movements in 3DPW and EMDB (1K frames from 3DPW and 2.5K frames from EMDB) for conducting the comparison. Note that 3DPW does not provide GT global trajectory annotations for evaluation. Results are shown in Tab. 5 and Fig. 11. We can conclude that our method achieved significant improvement in global translation estimation on EMDB (3DPW does not provide GT trans) and also produces slightly better pose estimation performance.
Figure 11: Comparison of 3D translation estimation in 3DPW.

### v2

@@ 4.2 Pose and Translation Optimization
Since we pay more attention to human body pose and global translation, we do not optimize the intrinsic parameter $K$, shape $\tilde{\boldsymbol{\beta}}$, and global rotation $\tilde{\boldsymbol{R}}$; instead use weak-perspective projection to estimate $K$ and adopt $\tilde{\boldsymbol{\beta}}$, $\tilde{\boldsymbol{R}}$, and initial pose $\tilde{\boldsymbol{\theta}}$ from CLIFF [31]. It is important to note that an image sequence corresponds to a specific individual. Therefore, we fix an average of $\tilde{\boldsymbol{\beta}}$ for the same person across the entire image sequence. To simplify the notation, we use $\boldsymbol{J}(\boldsymbol{\theta},\boldsymbol{T})$ to represent $\boldsymbol{J}(\boldsymbol{\theta},\tilde{\boldsymbol{\beta}},\tilde{\boldsymbol{R}},\boldsymbol{T})$.
@@
Contact joint loss. To recover the positional relationship between the human and the ground, we utilize the off-the-shelf depth estimation model ZoeDepth [2] to obtain a depth map for each RGB image and use the corresponding point cloud $\boldsymbol{P}$ to construct the 3D contact trace on the ground. Therefore, the spatial translation constraint term can be defined by penalizing the 3D distance between the SMPL foot joint $\boldsymbol{J}(\boldsymbol{\theta},\boldsymbol{T})_{foot}$ and the contact point cloud $\boldsymbol{P}(\mathcal{J}^{2d}_{foot})$:
@@
Foot consistency loss. Meanwhile, we assume that the foot joints’ position of two consecutive frames is consistently close during contact. At the same time, the jitter of 2D keypoints $\mathcal{J}^{2d}_{foot}$ will cause the corresponding contact joint on the ground to shift significantly. To solve the above problems, we define a temporal constraint term as
@@ 5.1 Comparison of Ground Truth Registration
We compare our RGBD-P fitting method with two existing RGBD methods: PROX [14] and LEMO [61]. Both of these methods consider human-scene contact in pose fitting. We employ the 3D depth error $E_{3d}$ to quantify the alignment between human models and depth observations. To evaluate the interaction between the model and the scene, we calculate the model contact by setting a unified threshold and compare the Mean Foot-Contact Error (MFCE), F1 score, and IOU with the ground truth contact. Tab. 2 demonstrates the effectiveness of the proposed method.
@@
Table 2: Evaluation of pose and shape registration on MMVP. M.: MFCE.
@@ 5.2 Comparison of Foot Contact Estimation
We apply precision, recall, F1 score, and IOU as evaluation metrics and conduct a quantitative comparison with BSTRO [18], BSTRO[FT], and DECO [49] on MMVP-test (Tab. 3). BSTRO[FT] represents the BSTRO fine-tuned on the MMVP training dataset. From Tab. 3, it is evident that our method achieves the best performance. It is worth mentioning that BSTRO and DECO, due to contact annotations in their training dataset (illustrated in Fig. 2), tend to predict feet as in contact for the majority of cases. As a result, their recall values are very high, while other metrics are undesirable. It is challenging to perceive the foot contact solely based on a single-frame image. By incorporating temporal motion information, we believe that’s why the proposed method surpasses the fine-tuned BSTRO.
To validate the generalization, we also compare our method with BSTRO[FT] on public datasets. Due to the inaccurate contact labels in other datasets, we only conduct qualitative comparisons (Fig. 5). It can be observed that the proposed method not only achieves more accurate contact predictions but also estimates reasonable foot pressure distributions. For more details, please refer to the supplementary material.
Figure 6: Qualitative comparison of 3D translation estimation in Side-Stepping (Left), Standing-Long-Jump (Middle) and Throwing-Medicine-Ball (Right) case. We estimate 3D translation stability and compare our results (Pink) with CLIFF [31] (Blue) and TRACE [48] (Green). The bottom row represents the acceleration of foot joints over time, which is used to measure the jitter of foot joints.
@@ 5.3 Pose and Translation Optimization Results
To assess the accuracy of our 3D pose estimation, we employ various metrics: MPJPE (Mean Per Joint Position Error), PMPJPE (Procrustes-aligned Mean Per Joint Position Error), PVE (Per Vertex Error), and PVE (Feet) which specifically focuses on the error of SMPL mesh vertices corresponding to the soles of the feet. For evaluating translation estimation, we utilize the offset distance of the pelvis point relative to the starting position, denoted as Traj.
Quantitative results. We conduct a comparative analysis of our method against CLIFF [31], TRACE [48], and SMPLer-X [5] on MMVP. As depicted in Tab. 5, our method exhibits satisfactory performance across all metrics. Notably, our method excels in Traj, indicating a significant improvement in acquiring more accurate and robust translation by leveraging foot contact for SMPL fitting.
Qualitative results. We compare 3D translation between our method, CLIFF [31], and TRACE [48] on MMVP (Fig. 6). In the Side-Stepping case, both CLIFF and TRACE exhibit drift in the depth direction, whereas our method demonstrates a consistent and stable translation in the depth direction. Similarly, in the Standing-Long-Jump and Throwing-Medicine-Ball case, the 3D translation estimated by CLIFF and TRACE manifests drift in the depth direction during the preparation stage. In contrast, our method avoids any depth-direction translation. At the bottom, we present the average absolute acceleration of foot joints which means the level of jitter. Our method achieves the smallest value, indicating superior foot stability.
5.4 Ablation Studies
To investigate the impacts of $L_{3d}$ and $L_{t}$, we conduct an ablation experiment on the MMVP dataset. As shown in Tab. 5, the omission of $L_{t}$ results in a marginal increase in all error metrics. This observation indicates that the temporal foot contact signals serve not only to improve the accuracy of translation but also to improve the accuracy of the pose. Without $L_{3d}$, MPJPE, PA-MPJPE, and PVE dropped slightly, but PVE (Feet) and Traj increased a lot. This is because when the 3D constraints of the feet and the ground are introduced, it will affect the 2D constraints to a certain extent and cause the accuracy of the pose to drop a little, but in exchange for the accuracy of the soles of the feet vertices and global translation. To attain enhanced translation, it is necessary to make a trade-off by sacrificing a certain degree of pose accuracy. Additionally, getting a better translation requires striking a balance between 2D and 3D constraints.
@@ 6 Conclusion
Limitations and future work: Our method is mainly designed for static viewpoints, so it remains challenging to handle moving camera configurations. The existing MMVP has relatively homogeneous scenes, making it difficult for current visual networks to achieve scene generalization on this dataset. In the future, we plan to collect data from various outdoor scenes to expand the dataset. Additionally, it would be an intriguing research endeavor to learn human body dynamics solely from pressure sequences.
Acknowledgements: This work was supported in part by the National Key R&D Program of China under Grant No.2022YFF0902201, the NSFC No.62171255, Xtep Science Lab, the Tsinghua University-Joint research and development project under Grant R24119F0 JCLFT-Phase 1, the “Ligiht Field Generic Technology Platform" (Z23111000290000) of Beijing Municipal Science and Technology Commission, the Aeronautical Science Fund under Grant 20230023051001, the Guoqiang Institute of Tsinghua University under Grant No.2021GQG0001.
@@ 7 Dataset Details
Synchronization: We uesed an Azure Kinect camera [^1] https://azure.microsoft.com/en-us/products/kinect-dk/ and Xsensor pressure insoles (HX 210-510) [^2] https://www.xsensor.com/solutions-and-platform/human-performance/gait-motion-insoles to capture the data. Due to this pressure insole’s lack of automatic synchronization support, we have designed a coarse-to-fine synchronization method. The pressing switch controls the bulb. When the switch is pressed with the foot, the bulb illuminates, and the insole detects the pressure simultaneously, achieving coarse synchronization. Manual fine synchronization is performed by observing the sudden changes in pressure on the insole, such as landing after a long jump.
Statistics: The dataset consists of a total of 16 subjects, with 11 males and 5 females. Key statistics (gender, age, height, and weight) of the subjects are shown in Fig. 7. Synchronized RGBD frames, pressure data, and registrations are more than 44k frames. The dataset consists of high-speed and large-range movements, including skipping rope, long jump, ball throwing, side stepping, running, and dancing.
@@ 8.1 Foot Contact Temporal Loss
We design four-foot planes derived from SMPL template mesh, as shown in Fig. 8. By projecting the foot vertices of the template SMPL model onto the ground, we acquire the anterior and posterior foot planes. The points on foot planes, which are associated with SMPL foot vertices, can also be controlled by SMPL parameters.
@@
8.2 RGBD-P Fitting Implementation Details
@@ 8.3 RGBD-P fitting Comparison Details
We compare our results with PROX  [^3] https://github.com/mohamedhassanmus/prox and LEMO  [^4] https://github.com/sanweiliti/LEMO on MMVP Dataset. All configs are set as their default values. We use four metrics to evaluate fitting performance: $E_{3d}$, Mean Foot-Contact Error (MFCE), F1 score, and IOU. For the last three metrics, given a fitting result, we set a fixed threshold to calculate the foot contact $C_{fitting}$ as POSA [15]. We calculate the F1 score and IOU between $C_{fitting}$ and the ground truth contact $C$. The Mean Foot-Contact Error (MFCE) is calculated as $\|C_{fitting}-C\|_{2}$.
Compared to PROX, our method considers body shape differences between adults and children, ensuring shape parameter consistency in one motion sequence. PROX sets a distance threshold between foot and ground, which may result in sudden motion. To avoid this, our method leverages dense foot-ground distance as loss function in optimization. Based on the fitting results of PROX, LEMO considers temporal motion changing and contact consistency, however, it neglects the alignment with depth data, which can lead to difficulties in accurately estimating global position.
@@ 9.1 FPP-Net
Network details. We follow PhysCap  [^5] https://github.com/soshishimada/PhysCap_demo_release and Jesse et al. [44] to construct our network. The details are shown in Fig. 9. The numbers next to "Conv.1d" represent the input dimension, output dimension, and kernel size. The numbers next to "Linear" represent the input dimension and output dimension. The outputs of FPP-Net are marked in purple. We train our network for 20 hours (900 epochs). The initial learning rate is $10^{-4}$. The optimizer is Adam.
Dataset split. We split 16 subjects in MMVP into 13 and 3 for training and testing, respectively.
@@ 9.3 Pose and Translation Optimization Comparison Details
Table 5: Evaluation of pose and translation estimation on 3DPW and EMDB.M.: MPJPE, PM. : PMPJPE, PF. : PVE (Feet).
Figure 11: Comparison of 3D translation estimation in 3DPW. From left to right: SMPLer-X (Purple), TRACE (Green), CLIFF (Blue), and our method (Pink).
9.4 More Results on In-the-Wild Datasets
We compared the results on in-the-wild datasets, 3DPW [53] and EMDB [24]. Our method is mainly designed for videos from static viewpoints, so it remains challenging to handle moving camera configurations. Therefore, we selected sequences with relatively static camera movements in 3DPW and EMDB (1K frames from 3DPW and 2.5K frames from EMDB) for conducting the comparison. Note that 3DPW does not provide GT global trajectory annotations for evaluation. As shown in Tab. 5 and Fig. 11, our method achieved significant improvement in global translation estimation on EMDB (3DPW does not provide GT trans) and also produces slightly better pose estimation results.